OpenStudy (anonymous):

Name the slope and y intercept of 2x + 5y =20

OpenStudy (saifoo.khan):

Make y the subject, \[y = \frac25x + \frac{20}{5}\] \[y = \frac25x + 4\]

OpenStudy (saifoo.khan):

Slope = co-efficent of x.

OpenStudy (anonymous):

slope: -2/5, y-int: (0,4)

OpenStudy (saifoo.khan):

wait, that's negative 2 over 5. i did a typo. sorry.

OpenStudy (anonymous):

How'd it become negative?

OpenStudy (anonymous):

solve the original equation for y. 5y = 20 - 2x so y = (20 - 2x)/5 or y = -2x/5 + 4